OpenStudy (anonymous):

find the slope, if it exist, of the line containing the pair of points (1,2) and (9,-6)

OpenStudy (dumbcow):

slope = rise/run slope = 8/-8 = -1

OpenStudy (lgbasallote):

slope = (y2-y1)/(x2-x1) so...-8/8 = -1

OpenStudy (anonymous):

so m=-8/8=-1? or os it 8/-8=-1

OpenStudy (anonymous):

the answer option is m=(blank)

OpenStudy (anonymous):

would the answer be m= -1